OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!! Write x^2 - 2x - 3 = 0 in the form (x - a)^2 = b, where a and b are integers. (x - 4)^2 = 3 (x - 3)^2 = 2 (x - 2)^2 = 1 (x - 1)^2 = 4

OpenStudy (anonymous):

@thomaster @dan815 @abb0t

OpenStudy (abb0t):

@dan815

OpenStudy (anonymous):

@ganeshie8 @zzr0ck3r

OpenStudy (zzr0ck3r):

@RosieF @abb0t @dan815

OpenStudy (anonymous):

y did u tag me on my own question @zzr0ck3r ?

OpenStudy (zzr0ck3r):

it seems like that was the thing to do

OpenStudy (zzr0ck3r):

\(x^2+bx+c=0 \implies (x+\frac{b}{2})^2=-c+(\frac{b}{2})^2\)

OpenStudy (zzr0ck3r):

\(b=-2, c=-3\)

OpenStudy (zzr0ck3r):

so b/2 = -1 and there is only one option

OpenStudy (anonymous):

so it's d?

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

thanks :)

OpenStudy (zzr0ck3r):

fo she z